Consumption function : C = 800 + 0.75Yd

Income : 1800

Investment : 500

Net export : 756

Government expenditure : 200

Tax : 200

Calculate the following value:

(i) Expenditure multiplier; (3 marks)

(ii) Total consumption; (3 marks)

(iii) Equilibrium level of income.


1
Expert's answer
2021-04-06T07:02:53-0400

(I) Expenditure multiplier

MPS= 0.75

MPC= 0.25

Expenditure multiplier=1.33

(ii) Total consumption

C= 800+ 0.75(1800)

C = 2,150

(iii)  Equilibrium level of income.

Y = C + I + G, where; Y is aggregate income,

C is consumption,

I is investment expenditure, and

G is government expenditure

therefore; 3456= 1800+ 500+ 756+200+200